OpenStudy (anonymous):

what is the solution system x+2y=-8 and x-y=13

OpenStudy (uri):

Multiply the second eq with 2

OpenStudy (uri):

x+2y=-8 x-y=13 (2) x+2y=-8 x-2y=26 ===== 2x=18 x=18/2 x=9 Now put the value of x in the original eq 1 x+2y=-8 9+2y=-8 Solve for y.